# Speaking Clearly: A Simplified Whisper-based Codec for Low-Bitrate Speech Coding

### v1

@@ 1 Introduction
In recent years, Speech Large Language Models (Speech LLMs) have garnered significant attention from the research community, demonstrating remarkable performance across a range of tasks [1, 2, 3, 4, 5]. The success of Speech LLMs is critically underpinned by a core component: the speech codec. This component serves as a crucial bridge, converting continuous audio signals into discrete tokens suitable for LLM modeling, thereby connecting raw audio with the model’s semantic understanding.
However, current speech codecs face an inherent conflict between the preservation of semantic content and acoustic fidelity, where optimizing for one typically degrades the other [6, 7]. This trade-off is particularly pronounced at low bitrates, where achieving high performance in both dimensions remains difficult. To mitigate this conflict, prevailing methods augment acoustic-centric codecs with external semantic supervision through various strategies. For instance, SpeechTokenizer [8] guides the first residual vector quantization (RVQ) layer through semantic distillation from HuBERT [9], while Mimi Codec in Moshi [6] employs split RVQ [10] with SSL model [11] supervision on one quantization branch. PAST [12] incorporates auxiliary phonetic tasks such as phoneme classification and ASR, and XY-Tokenizer [13] employs multi-task learning with LLM-based ASR supervision and multi-stage training. While effective, these methods typically rely on complex semantic supervision.
@@
  • Experimental results demonstrate that SimWhisper-Codec outperforms semantically-supervised codecs in semantic preservation at comparable low bitrates, while achieving high speaker similarity (0.83 SIM) and intelligibility (0.91 STOI), validating our method’s effectiveness.
@@ 2 EMPIRICAL ANALYSIS OF COMPONENTS HINDERING ACOUSTIC RECONSTRUCTION
We conduct an empirical analysis to identify which architectural components in Whisper encoders adversely affect acoustic reconstruction capabilities.
@@ 2.1 Component Analysis and Hypotheses
Absolute Positional Encodings. Absolute positional encodings assign fixed ”identity markers” to each temporal position in the sequence [19, 20]. We hypothesize that absolute positional encodings are detrimental to acoustic reconstruction because: (1) acoustic features should remain position-invariant—a phoneme /a/ should have identical representation regardless of temporal location; (2) speech contains repetitive structures that absolute encodings differentiate, hindering pattern recognition for reconstruction. These theoretical considerations motivate our experimental validation.
@@ 2.2 Validation Analysis
To validate our hypotheses, we conduct controlled analysis experiments using LJSpeech [21]. We extract frame-level hidden states from the final layer of each Whisper encoder variant, then condition identical HiFiGAN vocoders [22] on these features to assess reconstruction quality. This setup allows us to isolate the impact of specific architectural components on acoustic modeling capability while keeping the vocoder constant. Crucially, all encoder variants remain frozen during HiFiGAN training to serve as feature extractors.
We systematically examine the effect of removing each component individually, followed by their combined removal. The results validate our hypotheses: removing convolutional front-end nonlinearity yields substantial improvements with PESQ-NB increasing from 1.24 to 3.60 (+2.36), STOI from 0.82 to 0.96 (+0.14), and SIM from 0.81 to 0.86 (+0.05). Removing absolute positional encodings also confirms our hypothesis with PESQ-NB increasing from 1.24 to 2.95 (+1.71), STOI from 0.82 to 0.94 (+0.12), and SIM from 0.81 to 0.84 (+0.03).
Simultaneously removing both components yields the best reconstruction performance: PESQ-NB reaches 3.67 (+2.43), STOI achieves 0.97 (+0.15), and SIM attains 0.87 (+0.06). This demonstrates complementary hindering effects—the nonlinearity suppresses spectral details while positional encodings interfere with flexible attention patterns. Table 1 summarizes these results, with objective metrics defined in Section 4.2.
@@ 3.2 Model Architecture
Quantizer. We employ a Finite Scalar Quantization (FSQ) module [17], which mitigates codebook collapse and obviates the need for complex training machinery such as exponential moving averages and commitment losses required by traditional VQ methods.
Upsampler. The upsampler mirrors the downsampling architecture, employing nearest-neighbor upsampling combined with convolutional and residual blocks to restore the original temporal resolution. The module progressively expands the feature dimensionality back to the decoder’s input requirements.
@@ 3.3 Training Objective
where $D_{i}$ represents the $i$-th discriminator output from (MPD [22], MSD [26]), $N$ is the number of discriminators, and $G(z)$ is the generated audio. The generator adversarial loss is defined as:
@@ 4.1 Settings
Dataset and Training Details. We use the full training set of LibriSpeech [27] with 960 hours of speech data for training. The test-clean set with 2620 utterances is used for testing. All speech data are in 16 kHz with randomly cropped 4-second audio segments. Training is conducted on a single NVIDIA H100 GPU with batch size of 32 and gradient accumulation set to 3, resulting in an effective batch size of 96. The total training is performed for 500,000 steps using a single-stage approach. Both generator and discriminator employ AdamW optimization with $\beta_{1}=0.8$, $\beta_{2}=0.99$, and weight decay $0.01$. A cosine annealing learning rate schedule is used, declining from $1\times 10^{-4}$ to 0 with 5k warmup steps for both generator and discriminator.
@@
Baselines. We compare our codec against representative baselines at similar bitrates: EnCodec [7] (1.5 kbps), DAC-RVQ3 [29] (1.5 kbps), SpeechTokenizer [8] (1.0 kbps), Mimi-RVQ8 [6] (1.1 kbps), and XCodec2.0 [30] (0.8 kbps), all using official checkpoints.
@@ 4.2 Evaluation Metrics
SIM (speaker similarity). Cosine similarity is computed between speaker embeddings extracted from reference and reconstructed utterances using a WavLM-based speaker verification model [^1] https://github.com/microsoft/UniSpeech.
@@
4.3 Experimental Results
Table 2 presents a comprehensive comparison of SimWhisper-Codec against representative baselines. Our model effectively models both semantic and acoustic information simultaneously, validating our architectural simplification approach.
Semantic Preservation. On semantic evaluation, SimWhisper-Codec achieves the lowest Word Error Rate (WER) of 3.10, outperforming all other codecs. Notably, this performance is achieved without any semantic supervision, distinguishing it from baselines like XCodec2.0 (3.61 WER) and Mimi-RVQ8 (4.36 WER) that rely on such supervision. The results indicate that our architectural simplification does not compromise Whisper’s semantic alignment capabilities.
Acoustic Quality. In terms of acoustic reconstruction, SimWhisper-Codec delivers competitive results. It achieves a PESQ-NB of 2.98 and a STOI of 0.91, placing it among the top-performing models and significantly outperforming acoustic-only codecs like EnCodec and DAC-RVQ3. Furthermore, with a speaker similarity (SIM) score of 0.83, it demonstrates good preservation of speaker identity. These results show that the simplified Whisper encoder can effectively extract acoustic features even at low bitrates of 1.1 kbps.
@@ 4.3 Experimental Results
Ground Truth 2.53
EnCodec [7] 1.5k 75 No 45.49 0.60 0.85 1.94 1.56
DAC-RVQ3 [29] 1.5k 75 No 41.67 0.45 0.76 1.82 1.43
SpeechTokenizer [8] 1.0k 50 Yes 5.92 0.37 0.70 1.42 1.15
Mimi-RVQ8 [6] 1.1k 12.5 Yes 4.36 0.73 0.90 2.62 2.13
XCodec2.0 [30] 0.8k 50 Yes 3.61 0.82 0.91 2.95 2.32
SimWhisper-Codec (ours) 1.1k 12.5 No 3.10 0.83 0.91 2.98 2.36
@@ 4.4 Ablation Study
Ours: remove both components 3.10 0.83 0.91 2.98 2.36
Table 3 demonstrates that both architectural modifications contribute to performance improvements. Removing GELU activations from the convolutional front-end has a more substantial impact on semantic preservation (WER: 5.95→3.74), while removing positional encodings provides moderate gains across all metrics. The combination of both modifications yields the best results, achieving a WER of 3.10, SIM of 0.83, and PESQ-WB of 2.36, validating our architectural design choices.
@@ 5 Conclusion
We presented SimWhisper-Codec, a low-bitrate speech codec that mitigates the semantic-acoustic conflict through architectural simplification rather than complex supervision. We removed convolutional front-end nonlinearity and absolute positional encodings from the frozen Whisper encoder. This architectural simplification maintains competitive acoustic quality while achieving superior semantic preservation. Results demonstrate that such simplification of Whisper can be more effective than semantic supervision approaches for speech codec design.

### v2

@@ 1 Introduction
In recent years, Speech Large Language Models (Speech LLMs) have garnered significant attention from the research community, demonstrating remarkable performance across a range of tasks [1, 2, 3, 4]. The success of Speech LLMs is critically underpinned by a core component: the speech codec. This component serves as a crucial bridge, converting continuous audio signals into discrete tokens suitable for LLM modeling[5], thereby connecting raw audio with the model’s semantic understanding.
However, current speech codecs face an inherent conflict between the preservation of semantic content and acoustic fidelity, as optimizing for one typically degrades the other [6, 7]. This trade-off is particularly pronounced at low bitrates, where achieving high performance in both dimensions remains difficult. To mitigate this conflict, prevailing methods augment acoustic-centric codecs with external semantic supervision through various strategies. For instance, SpeechTokenizer [8] guides the first residual vector quantization (RVQ) layer through semantic distillation from HuBERT [9], while Mimi Codec in Moshi [6] employs split RVQ [10] with SSL model [11] supervision on one quantization branch. PAST [12] incorporates auxiliary phonetic tasks such as phoneme classification and ASR, and XY-Tokenizer [13] employs multi-task learning with LLM-based ASR supervision and multi-stage training. While effective, these methods typically rely on complex semantic supervision.
@@
  • Experimental results demonstrate that SimWhisper-Codec achieves state-of-the-art acoustic quality (3.29 PESQ-NB, 2.72 PESQ-WB) and competitive semantic preservation (2.75 WER) at similarly low bitrates without external supervision, validating the effectiveness of our approach.
@@ 2 EMPIRICAL ANALYSIS OF COMPONENTS HINDERING ACOUSTIC RECONSTRUCTION
We conduct an empirical analysis to identify which architectural components in multilingual Whisper encoders adversely affect acoustic reconstruction capabilities.
@@ 2.1 Component Analysis and Hypotheses
Absolute Positional Encodings. Absolute positional encodings assign fixed “identity markers” to each temporal position in the sequence [19, 20]. We hypothesize that absolute positional encodings are detrimental to acoustic reconstruction because: (1) acoustic features should remain position-invariant—a phoneme /a/ should have identical representation regardless of temporal location; (2) speech contains repetitive structures that absolute encodings differentiate, hindering pattern recognition for reconstruction. These theoretical considerations motivate our experimental validation.
@@ 2.2 Validation Analysis
To validate our hypotheses, we conduct controlled analysis experiments using LJSpeech [^1] https://keithito.com/LJ-Speech-Dataset/. We extract frame-level hidden states from the final layer of each Whisper encoder variant, then condition identical HiFiGAN vocoders [21] on these features to assess reconstruction quality. This setup allows us to isolate the impact of specific architectural components on acoustic modeling capability while keeping the vocoder constant. Crucially, all encoder variants remain frozen during HiFiGAN training to serve as feature extractors [22].
We examine the effect of removing each component individually, then their combined removal. The results confirm our hypotheses: removing convolutional front-end nonlinearity yields substantial improvements with PESQ-NB increasing from 1.24 to 3.60 (+2.36), STOI from 0.82 to 0.96 (+0.14), and SIM from 0.81 to 0.86 (+0.05). Removing absolute positional encodings also confirms this with PESQ-NB increasing from 1.24 to 2.95 (+1.71), STOI from 0.82 to 0.94 (+0.12), and SIM from 0.81 to 0.84 (+0.03).
Simultaneously removing both components yields the best reconstruction performance: PESQ-NB reaches 3.67 (+2.43), STOI achieves 0.97 (+0.15), and SIM attains 0.87 (+0.06). This demonstrates complementary hindering effects—the nonlinearity suppresses spectral details while positional encodings interfere with flexible attention patterns. Table 1 summarizes these results, with metrics defined in Section 4.2. We validate that these improvements generalize across the entire Whisper model family, from tiny to large-v3, demonstrating the robustness of our findings.
@@ 3.2 Model Architecture
Quantizer. We employ a Finite Scalar Quantization (FSQ) module [17], which mitigates codebook collapse and obviates the need for complex training machinery such as exponential moving averages and commitment losses required by traditional VQ [24] methods.
Upsampler. The upsampler first reconstructs features through residual blocks with dilated convolutions and Snake activations. It then expands the channel dimension and unstacks the features to restore the original temporal resolution and dimensionality for the decoder.
@@ 3.3 Training Objective
where $D_{i}$ represents the $i$-th discriminator output from (MPD [21], MS-STFTD [7]), $N$ is the number of discriminators, and $G(z)$ is the generated audio. The generator adversarial loss is defined as:
@@ 4.1 Settings
Dataset and Training Details. We use the full training set of LibriSpeech [27] with 960 hours of speech data for training. The test-clean set with 2620 utterances is used for testing. All speech data are in 16 kHz with randomly cropped 2-second audio segments. Training is conducted on a single NVIDIA H100 GPU with batch size of 64 and gradient accumulation set to 1, resulting in an effective batch size of 64. The total training is performed for 1000,000 steps using a single-stage approach. Both generator and discriminator employ AdamW optimization with $\beta_{1}=0.8$, $\beta_{2}=0.99$, and weight decay $0.01$. A cosine annealing learning rate schedule is used, declining from $1\times 10^{-4}$ to 0 with 10k warmup steps for both components.
@@
Baselines. We compare our codec against representative baselines at similar bitrates: EnCodec [7] (1.5 kbps), DAC-RVQ3 [29] (1.5 kbps), SpeechTokenizer [8] (1.0 kbps), Mimi-RVQ8 [6] (1.1 kbps), BigCodec(1.04 kbps), XCodec2.0 [30] (0.8 kbps) and XY-Tokenizer [13] (1.0 kbps), all using official checkpoints.
@@ 4.2 Evaluation Metrics
SIM (speaker similarity). Cosine similarity is computed between speaker embeddings extracted from reference and reconstructed utterances using a WavLM-based speaker verification model [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification.
@@
Ground Truth 2.16 1.00 1.00 4.55 4.64
EnCodec [7] 1.5k 75 No 5.62 0.60 0.85 1.94 1.56
DAC-RVQ3 [29] 1.5k 75 No 7.80 0.45 0.76 1.82 1.43
SpeechTokenizer [8] 1.0k 50 Yes 4.21 0.37 0.70 1.42 1.15
BigCodec [33] 1.04k 80 No 2.92 0.84 0.93 3.26 2.68
Mimi-RVQ8 [6] 1.1k 12.5 Yes 3.24 0.73 0.90 2.79 2.24
XCodec2.0 [30] 0.8k 50 Yes 2.61 0.82 0.92 3.04 2.43
XY-Tokenizer [13] 1.0k 12.5 Yes 2.46 0.85 0.92 3.10 2.50
SimWhisper-Codec (ours) 1.1k 12.5 No 2.75 0.83 0.93 3.29 2.72
4.3 Experimental Results
Table 2 presents a comprehensive comparison of SimWhisper-Codec against representative baselines. Our approach effectively models both semantic and acoustic information simultaneously, validating our architectural simplification.
Semantic Preservation. SimWhisper-Codec achieves a competitive WER of 2.75, comparable to strong supervised baselines like XY-Tokenizer (2.46) and XCodec2.0 (2.61). Notably, this is achieved without semantic supervision, unlike these baselines which rely on complex distillation or multi-task learning. This suggests that our architectural simplification effectively preserves Whisper’s semantic alignment capabilities.
Acoustic Quality. For acoustic reconstruction, SimWhisper-Codec delivers state-of-the-art performance. It achieves 3.29 PESQ-NB and 2.72 PESQ-WB. These scores surpass the leading acoustic-only baseline, BigCodec (3.26 and 2.68). It also matches BigCodec’s high intelligibility (0.93 STOI). Furthermore, with a speaker similarity (SIM) score of 0.83, it demonstrates good preservation of speaker identity. These results validate our hypothesis: targeted architectural simplification unlocks Whisper’s acoustic modeling potential.
@@ 4.4 Ablation Study
Ours: remove both components 2.75 0.83 0.93 3.29 2.72
Table 3 demonstrates that both architectural modifications contribute to performance improvements. Removing GELU activations from the convolutional front-end has a more substantial impact on semantic preservation (WER: 5.95→3.74), while removing positional encodings provides moderate gains across all metrics. The combination of both modifications yields the best results, achieving a WER of 2.75, SIM of 0.83, and PESQ-WB of 2.72, validating our architectural design choices.
@@ 5 Conclusion
We presented SimWhisper-Codec, a low-bitrate speech codec that mitigates the semantic-acoustic conflict through architectural simplification rather than complex supervision. We removed convolutional front-end nonlinearity and absolute positional encodings from the frozen Whisper encoder. This architectural simplification achieves superior acoustic quality while maintaining competitive semantic preservation. Results demonstrate that such simplification of Whisper can be more effective than semantic supervision approaches for speech codec design.